solve 6(4 - 5x)= 2 - 2(3 -x)show work step by step.
Answer by stanbon(75887)

solve 6(4 - 5x)= 2 - 2(3 -x)
---
24 - 30x = 2 - 6 + 2x
24-30x = -4 + 2x
32x = 28
x = 7/8
=================
Cheers,
Stan H.
=================